OpenStudy (anonymous):

What is the sum of all even integers between 99 and 301?

OpenStudy (anonymous):

first term = 100 , last term= 300 , common difference =2 no. of terms = (300-100)/2 = 100 sum of all even numbers between 99 and 301 = 100/2 * (100 + 300) = 50*400=20000

OpenStudy (anonymous):

But isn't 200 only counted once? So would it be 19800?

OpenStudy (anonymous):

yeah..20000

OpenStudy (anonymous):

Its 20000 im pretty positive.

OpenStudy (anonymous):

Ill check this

OpenStudy (anonymous):

Im convinced that its actually 20200

OpenStudy (midhun.madhu1987):